OpenStudy (anonymous):

Use the product of a sum and a difference to evaluate the expression. 43x37=

OpenStudy (mayankdevnani):

(40+3)(40-3)

OpenStudy (mayankdevnani):

40^2-3^2 1600-9=1591

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{(a+b)(a-b)=a^2-b^2}}\]\[\Huge{\color{green}{43 \times 37=(40+3)(40-3)}}\]\[\Huge{\color{blue}{(40)^2-(3)^2}}\]

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{purple}{1600-9=1591.}}\]